# Computing the Intrinsic Delaunay Triangulation of a Closed Polyhedral Surface

### v1

@@ I.3 Voronoi diagram and Delaunay tessellation
Proof J.28.
Consider $i\in[k]$, and assume that the point $z_{i}$ has just been sent by Alice. We must return to Alice. The crux of the matter is to have maintained at this point the list of tuples $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ over $j\in[i-1]$, ordered by the position of $\text{Vor}(z_{j},Z_{i-1})\cap I$ along $I$ (for some direction of $I$, and resolving any ambiguity arbitrarily). Now we can use the list to answer Alice, and update the list, as follows. Given a tuple $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ we can determine in constant time whether $j\in U_{i}$ by checking whether there is a point of $\text{Vor}(j,Z_{i-1})\cap I$ that is strictly closer to $z_{i}$ than to $z_{j}$. If $j\notin U_{i}$, then either all the tuples $(j^{\prime},\text{Vor}(z_{j^{\prime}},Z_{i-1})\cap I)$ before $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ in the list are such that $j^{\prime}\notin U_{i}$, or all the tuples after $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ are like that, and we can find out which case it is in constant time. So we can list by dichotomy the $k^{\prime}\geq 0$ tuples $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ such that $j\in U_{i}$ in $O(k^{\prime}+\log k)$ time. For each such tuple $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$, we derive $\text{Vor}(z_{j},Z_{i})\cap I$ from $\text{Vor}(z_{j},Z_{i-1})$, $z_{j}$, and $z_{i}$ in constant time. In the end we compute $\text{Vor}(z_{i},Z_{i})\cap I$ in $O(\log k)$ time by finding by dichotomy the first and last tuples $(j,\text{Vor}(z_{j},Z_{i-1})$ such that $\text{Vor}(z_{j},Z_{i-1})\cap I$ contains a point that is at least as close to $z_{i}$ than to $z_{j}$, if any. This way we can return to Alice, and update the list of tuples, in $O(k^{\prime}+\log k)$ total time. Lemma J.25 concludes.
@@
Finally, Lemma J.27 also provides us with the set $\text{Vor}(x,X_{\Delta}\cup\{x\})\cap s$. If this set is not empty, and if $\lambda(x)\notin X_{\Delta^{\prime}}$, then we consider the distance $t$ between $x$ and $\text{Vor}(x,X_{\Delta}\cup\{x\})\cap s$, and we create the event $(t,s^{\prime},\Delta^{\prime},\lambda(x))$, in $O(\log k)$ time.
@@
Theorem 1.1 has two interesting applications. First, recall that Delaunay triangulations have bounded happiness [portalgons, Section 4]. Combined with Theorem 1.1 we obtain:
@@
Proof K.2 (Proof of Corollary K.1).
Apply Theorem 1.1 to compute the portalgon $T^{\prime}$ of the Delaunay tessellation of $\mathcal{S}(T)$ in $O(n^{3}\log^{2}(n)\cdot\log^{4}(r))$ time. Some polygons of $T^{\prime}$ may not be triangles. Cut the polygons of $T^{\prime}$ that are not triangles (if any) along vertex-to-vertex arcs to obtain a triangular portalgon $T^{\prime\prime}$. Then $T^{\prime\prime}$ is the portalgon of a Delaunay triangulation of $\mathcal{S}(T)$, so $T^{\prime\prime}$ has bounded happiness by the result of Löffler, Ophelders, Silveira, and Staals [portalgons, Section 4]. Moreover the vertex set of its 1-skeleton $T^{\prime\prime 1}$ is exactly the set of singularities of $\mathcal{S}(T)$, except if $\mathcal{S}(T)$ is a flat torus in which case $T^{\prime\prime 1}$ has exactly one vertex, so in any case $T^{\prime\prime}$ has $O(n)$ triangles.
On the portalgon $T^{\prime}$ returned by Corollary K.1 the single-source shortest path algorithm of Löffler, Ophelders, Silveira, and Staals [portalgons, Section 3] would run in time $O(n^{2}\log^{O(1)}(n))$, so that:
{observation}
@@
Let us prove the claim. We consider the sides of the polygons of $\mathcal{T}$ and $\mathcal{T}^{\prime}$. There are $O(n)$ such sides. Fix a side $s$ of a polygon of $\mathcal{T}$. For every side $s^{\prime}$ of a polygon of $\mathcal{T}^{\prime}$, determine in $O(n)$ time whether there exists a one-to-one correspondence $\varphi$ from the sides of the polygons of $\mathcal{T}$ to the sides of the polygons of $\mathcal{T}^{\prime}$ that maps $s$ to $s^{\prime}$, the boundary closed walks of the polygons of $\mathcal{T}$ to the boundary closed walks of the polygons of $\mathcal{T}^{\prime}$, and the matching of $\mathcal{T}$ to the matching of $\mathcal{T}^{\prime}$. If $\varphi$ exists then $\varphi$ is unique since $\mathcal{S}(\mathcal{T})$ and $\mathcal{S}(\mathcal{T}^{\prime})$ are connected: construct $\varphi$ in $O(n)$ time. Then determine in $O(n)$ time if for every polygon $P$ of $\mathcal{T}$ there is an orientation-preserving isometry $\tau_{P}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ such that $\varphi(s)=\tau_{P}(s)$ for every side $s$ of $P$. In which case return correctly that $\mathcal{T}$ and $\mathcal{T}^{\prime}$ are equal. In the end, if every polygon side $s^{\prime}$ of $\mathcal{T}^{\prime}$ has been looped upon, and if no equality has been found, return correctly that $\mathcal{T}$ and $\mathcal{T}^{\prime}$ are distinct. This proves the claim, and the corollary.

### v2

@@ I.3 Voronoi diagram and Delaunay tessellation
Proof.
Consider $i\in[k]$, and assume that the point $z_{i}$ has just been sent by Alice. We must return to Alice. The crux of the matter is to have maintained at this point the list of tuples $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ over $j\in[i-1]$, ordered by the position of $\text{Vor}(z_{j},Z_{i-1})\cap I$ along $I$ (for some direction of $I$, and resolving any ambiguity arbitrarily). Now we can use the list to answer Alice, and update the list, as follows. Given a tuple $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ we can determine in constant time whether $j\in U_{i}$ by checking whether there is a point of $\text{Vor}(j,Z_{i-1})\cap I$ that is strictly closer to $z_{i}$ than to $z_{j}$. If $j\notin U_{i}$, then either all the tuples $(j^{\prime},\text{Vor}(z_{j^{\prime}},Z_{i-1})\cap I)$ before $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ in the list are such that $j^{\prime}\notin U_{i}$, or all the tuples after $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ are like that, and we can find out which case it is in constant time. So we can list by dichotomy the $k^{\prime}\geq 0$ tuples $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$ such that $j\in U_{i}$ in $O(k^{\prime}+\log k)$ time. For each such tuple $(j,\text{Vor}(z_{j},Z_{i-1})\cap I)$, we derive $\text{Vor}(z_{j},Z_{i})\cap I$ from $\text{Vor}(z_{j},Z_{i-1})$, $z_{j}$, and $z_{i}$ in constant time. In the end we compute $\text{Vor}(z_{i},Z_{i})\cap I$ in $O(\log k)$ time by finding by dichotomy the first and last tuples $(j,\text{Vor}(z_{j},Z_{i-1})$ such that $\text{Vor}(z_{j},Z_{i-1})\cap I$ contains a point that is at least as close to $z_{i}$ than to $z_{j}$, if any. This way we can return to Alice, and update the list of tuples, in $O(k^{\prime}+\log k)$ total time. Lemma J.13 concludes. ∎
@@
Finally, Lemma J.14 also provides us with the set $\text{Vor}(x,X_{\Delta}\cup\{x\})\cap s$. If this set is not empty, and if $\lambda(x)\notin X_{\Delta^{\prime}}$, then we consider the distance $t$ between $x$ and $\text{Vor}(x,X_{\Delta}\cup\{x\})\cap s$, and we create the event $(t,s^{\prime},\Delta^{\prime},\lambda(x))$, in $O(\log k)$ time. ∎
@@
Theorem 1.1 has two interesting applications. First, recall that Delaunay triangulations have bounded happiness [22, Section 4]. Combined with Theorem 1.1 we obtain:
@@
Proof of Corollary K.1.
Apply Theorem 1.1 to compute the portalgon $T^{\prime}$ of the Delaunay tessellation of $\mathcal{S}(T)$ in $O(n^{3}\log^{2}(n)\cdot\log^{4}(r))$ time. Some polygons of $T^{\prime}$ may not be triangles. Cut the polygons of $T^{\prime}$ that are not triangles (if any) along vertex-to-vertex arcs to obtain a triangular portalgon $T^{\prime\prime}$. Then $T^{\prime\prime}$ is the portalgon of a Delaunay triangulation of $\mathcal{S}(T)$, so $T^{\prime\prime}$ has bounded happiness by the result of Löffler, Ophelders, Silveira, and Staals [22, Section 4]. Moreover the vertex set of its 1-skeleton $T^{\prime\prime 1}$ is exactly the set of singularities of $\mathcal{S}(T)$, except if $\mathcal{S}(T)$ is a flat torus in which case $T^{\prime\prime 1}$ has exactly one vertex, so in any case $T^{\prime\prime}$ has $O(n)$ triangles. ∎
On the portalgon $T^{\prime}$ returned by Corollary K.1 the single-source shortest path algorithm of Löffler, Ophelders, Silveira, and Staals [22, Section 3] would run in time $O(n^{2}\log^{O(1)}(n))$, so that:
Observation 1.
@@
Let us prove the claim. We consider the sides of the polygons of $\mathcal{T}$ and $\mathcal{T}^{\prime}$. There are $O(n)$ such sides. Fix a side $s$ of a polygon of $\mathcal{T}$. For every side $s^{\prime}$ of a polygon of $\mathcal{T}^{\prime}$, determine in $O(n)$ time whether there exists a one-to-one correspondence $\varphi$ from the sides of the polygons of $\mathcal{T}$ to the sides of the polygons of $\mathcal{T}^{\prime}$ that maps $s$ to $s^{\prime}$, the boundary closed walks of the polygons of $\mathcal{T}$ to the boundary closed walks of the polygons of $\mathcal{T}^{\prime}$, and the matching of $\mathcal{T}$ to the matching of $\mathcal{T}^{\prime}$. If $\varphi$ exists then $\varphi$ is unique since $\mathcal{S}(\mathcal{T})$ and $\mathcal{S}(\mathcal{T}^{\prime})$ are connected: construct $\varphi$ in $O(n)$ time. Then determine in $O(n)$ time if for every polygon $P$ of $\mathcal{T}$ there is an orientation-preserving isometry $\tau_{P}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ such that $\varphi(s)=\tau_{P}(s)$ for every side $s$ of $P$. In which case return correctly that $\mathcal{T}$ and $\mathcal{T}^{\prime}$ are equal. In the end, if every polygon side $s^{\prime}$ of $\mathcal{T}^{\prime}$ has been looped upon, and if no equality has been found, return correctly that $\mathcal{T}$ and $\mathcal{T}^{\prime}$ are distinct. This proves the claim, and the corollary. ∎